OpenStudy (anonymous):

in the xy-plane which of the following points lies inside the circular region of radius 3 centered at (1,-2)

OpenStudy (anonymous):

any choises?

OpenStudy (anonymous):

A- (-1,1) B (1,2) C (0,-5) D (3,0)

OpenStudy (anonymous):

well let me think

OpenStudy (anonymous):

well one more minute

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

D i think

OpenStudy (mayankdevnani):

can you write equation of circle using given information ? @Nicomia1023

OpenStudy (anonymous):

(x-1)^2+(y+2)^2=9

OpenStudy (mayankdevnani):

good !!

OpenStudy (anonymous):

well in the region of this circle only D lies inside

OpenStudy (mayankdevnani):

after simplifying it, we get \[\large \bf x^2+y^2-2x+4y-4=0\]

OpenStudy (anonymous):

yes

OpenStudy (mayankdevnani):

now plug all the values in the equation ,if the value come out which is less than 0,then that point will lie inside the circular region and if the value come out which is equal to 0,then that point will lie on the circular region and if the value come out is greater than 0,then that point will lie outside the circle.

OpenStudy (anonymous):

so whats your answer?

OpenStudy (anonymous):

@Nicomia1023 can you give me a medal cuz i already gave a medal to @mayankdevnani

OpenStudy (mayankdevnani):

plug value of option D i.e (3,0) we get \[\large \bf 9+0-6+0-4=-1\] we can see that value is less than 0,so option D is correct answer :)

OpenStudy (mayankdevnani):

hope you understand. @Nicomia1023

OpenStudy (anonymous):

i understand a little

OpenStudy (mayankdevnani):

and good job !@ttb123456789 but try to explain the ask more clear about solving a question

OpenStudy (mayankdevnani):

@ttb123456789

OpenStudy (anonymous):

ok

OpenStudy (mayankdevnani):

good !! @Nicomia1023

OpenStudy (mayankdevnani):

*asker in previous to previous comment

OpenStudy (mayankdevnani):

you can also give MEDAL by clicking BEST RESPONSE flashing right side to my and @ttb123456789 's comment `But choose that user which helps you a lot`

OpenStudy (mayankdevnani):

@Nicomia1023

OpenStudy (anonymous):

thank you